OpenStudy (watson):

integrate (tan(x)+1)/(tan(x)-1)..help me solve this please..

OpenStudy (lgbasallote):

hmm wait..it seems it'swritten \[\frac{\tan x + 1}{\tan x - 1}\] sorry misread

OpenStudy (watson):

how to integrate that??

OpenStudy (anonymous):

\[\large\int\limits\frac{\frac{sinx}{cosx}+1}{\frac{sinx}{cosx}-1} dx \] Weierstrass substitution might work..

OpenStudy (experimentx):

try multiplying by conjugate (what it is called??) of denominator and ... change into double angle

OpenStudy (anonymous):

wont weisestrass substitution work? http://en.wikipedia.org/wiki/Weierstrass_substitution

OpenStudy (watson):

if i take u is tan(x)+1 it won't work

OpenStudy (anonymous):

indeed it wont work. try manipulating it with trig identities; changding tanx = sinx/cosx like i suggested.

OpenStudy (watson):

ok i will try..thanks :)

OpenStudy (experimentx):

this should yield to something like \[ \frac{1 + \sin 2x}{-\cos 2x} \]

OpenStudy (watson):

is that by using trigonometry identities?